OpenStudy (anonymous):

6. Find the third term in this sequence: tn=2n/5 (Points : 1)

OpenStudy (anonymous):

put n= 3 t3=2*3/5 =6/5

OpenStudy (mathstudent55):

\(t_3 = \dfrac{2(3)}{5} \)